OpenStudy (anonymous):

Angle A in Quadrant III and 0°≤A<360°, cosA= -2/5 cos A/2= a. √7/10 b. √3/10 c. -√3/10 d. -√7/10

OpenStudy (ash2326):

It's given that A is in third quadrant, so A would lie between 180 and 270 degrees \[180\le A \le 270\] Can you tell the range of A/2 ?

OpenStudy (anonymous):

Oh o.k. Wait, the range of A/2?

OpenStudy (ash2326):

yeah, like A is from 180 to 270 what would be for A/2?

OpenStudy (anonymous):

90 to 135?

OpenStudy (ash2326):

good, so what's the sign of cosine in this range?

OpenStudy (ash2326):

@mathsux4real what's the sign of cos in II quadrant?

OpenStudy (anonymous):

cos(θ) < 0 ?

OpenStudy (ash2326):

good:) so here we would have cos A/2 <0 do yo know the half angle formula for cosine?

OpenStudy (anonymous):

Can you give it to me? I don't know it right off the bat

OpenStudy (ash2326):

it's given as \[\Large 2 \cos^2 \frac A 2=1+ \cos A\]

OpenStudy (ash2326):

find cos A/2 and remember that it's negative

OpenStudy (anonymous):

I can't seem to match it up.

OpenStudy (ash2326):

put cos A=-2/5 I think you'd get it. Try again :)

OpenStudy (anonymous):

So 0.6

OpenStudy (ash2326):

it's \[2\times \cos^2 \frac A 2=1+\cos A\] you have to divide by 2 and then take root

OpenStudy (anonymous):

I came up with either 1.825 or 1.35 :/

OpenStudy (anonymous):

@dta255 any thoughts?

OpenStudy (ash2326):

\[2\times \cos^2 \frac A 2=1-\frac 25\] \[2\times \cos^2 \frac A 2=\frac 35\] divide both sides by 2, we get \[ \cos^2 \frac A 2=\frac 3{10}\] Taking root both sides, we get \[ \cos \frac A 2=\pm \sqrt {\frac {3}{10}}\] Now which sign should we take

OpenStudy (anonymous):

+

OpenStudy (ash2326):

@mathsux4real here \[\frac A 2\] lies in the second quadrant, then what will be the sign?

OpenStudy (anonymous):

-

OpenStudy (ash2326):

yes:) do you understand this?

OpenStudy (anonymous):

Yes. I looked it up to understand fully. Thank you very much ash

OpenStudy (ash2326):

welcome:D